### v3

# Descriptions of Cantor Sets: A Set‑Theoretic Survey and Open Problems

Abstract
This survey synthesizes the principal descriptive set-theoretic perspectives on deterministic Cantor sets on the real line and charts directions for future study. After recounting their historical genesis and compiling an up-to-date taxonomy, we review the Borel hierarchy and four hierarchically ordered representations—general, nested, iterated-function-system (IFS), and $q$-ary expansion—presented from the most general to the most specific set-theoretic description of deterministic Cantor sets. We then present explicit and recursive descriptions for two thin families of measure-zero Cantor sets and an augmented “tick" family of positive measure, respectively, showing that the classical middle-third set lies in the intersection of all three families of after-mentioned Cantor sets. The survey closes by isolating several open problems in four directions, aiming to provide mathematicians with a coherent platform for further descriptive set-theoretic investigations into Cantor-type sets on the real line.
Mathematics Subject Classification (2020): 03E05, 28A80, 54A05
Keywords Cantor set, Borel operations, descriptive formula, set theory, fractal geometry
“A set is a Many that allows itself to be thought of as a One.”
— Georg Cantor (1845–1918)
1 Introduction
The Introduction is structured as follows. Section 1.1 reviews the historical development of Cantor sets, while Section 1.2 presents an up-to-date taxonomy of their principal variants. Section 1.3 sets out the motivation for undertaking a set-theoretic survey of these constructions, and Section 1.4 closes the Introduction with a synopsis of the paper’s overall organization.
1.1 Historical Background
The conceptual roots of the Cantor set lie in Georg Cantor’s 1872 introduction of the derived set, which isolated the limit points of a set and laid the groundwork for perfect, nowhere‐dense subsets of $\mathbb{R}$.  Shortly thereafter, H. J. S. Smith’s 1875 note provided the first printed construction of a “Cantor‐type” set by iteratively excising subintervals of $[0,1]$, thereby demonstrating that a closed set can be uncountable yet possess arbitrarily small outer content, although the significance of his example went largely unnoticed at the time.  Cantor returned to the theme in his six–part memoir Über unendliche lineare Punktmannichfaltigkeiten (1879–1884); Part V (written October 1882, published 1883) contains the now‐classical 3-ary or ternary expansion set (later called the middle-third Cantor set), presented as a perfect set that is nowhere dense.  In a letter dated November 1883 Cantor extends this construction of the Cantor set by defining the Cantor function, a continuous, non‐decreasing map that is constant on the removed intervals and thus furnishes a counter-example to naïve forms of the Fundamental Theorem of Calculus.  His 1884 paper De la puissance des ensembles parfaits de points in Acta Mathematica analyzed the cardinalities of perfect sets, cementing the role of the ternary Cantor set as both a set-theoretic and analytic paradigm. These milestones—Smith’s overlooked precursor, Cantor’s derived-set framework, the explicit ternary construction, and the analytic Cantor function—together mark the emergence of what is now simply known as the Cantor set [1].
1.2 Taxonomy of Cantor Sets
Various classes and approaches to the construction of Cantor sets have been extensively studied, each generating distinct subclasses and mathematical structures. As of the date of this survey (15 June 2025), mathematicians have identified numerous specific types of Cantor sets, including Affine, Autonomous, Balanced, Bunched, Central, Cookie-Cutter, Deterministic, Discrete, Distorted, Dusts, Fat, Generalized, Genus-g, Geometric, Hairy, High Dimensional, Homogeneous, Hyperbolic, Iteration Function System (IFS) type, Invariant, Kinetic, Linear, Minimal, Multi-model, Nearly-affine, Non autonomous, Non-homogeneous, Non-archmedian (p-adic), Non-central, Non-Symmetric, Nonstandard, p-Adic Heisenberg, Random, Regular, Scrawny, Self-Similar, Semi-bounded, Smooth, Sticky, Superior, Smith-Volterra-Cantor(SVC) type, Symbolic, Symmetric, Tame, Thick, Thin, Twofold, Ultrametric, Uniform, Universal, Visible, and Wild. Comprehensive references for these subclasses can be found in [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45]. In this paper, however, our primary focus is on deterministic Cantor sets defined explicitly on the real line, and we direct the interested reader to relevant literature for more general or abstract variants.
1.3 Motivation and Scope
Since Georg Cantor’s original construction in the 1880s, the Cantor set has become a touchstone throughout mathematics and beyond. On the real line it provides one of the simplest explicit exemplars of a fractal: it is nowhere dense, totally disconnected, perfect, and uncountable, yet its Lebesgue measure can be tuned from zero (the classical middle-third set) to any prescribed value in $(0,1).$ Such extremal combinations of topological and measure-theoretic properties make Cantor sets a natural testing ground for “pathological" phenomena—most famously the Cantor–Lebesgue function, a continuous, non-decreasing map that is constant on a set of full measure. In constructive and descriptive settings the Cantor set reappears as a paradigmatic building block: for example, basic Cantor-type function blocks span 5/28 of the canonical representatives in the function space $F(R,R),$ [46] and they appear to prove the existence of aleph-two deterministic as well as aleph-two random fractals on the real line [47, 48]. Outside pure mathematics, Cantor-type patterns have been cited in art—early Egyptian column motifs [49, 50], in modeling financial markets [51] —and in nature, such as ring-like stratification of Saturn when idealized as repeated circular products [52]. Despite this wide relevance, no unified survey has yet cataloged the diverse set-theoretic descriptions of Cantor sets, nor identified which directions are fully developed and which remain open. The present paper closes that gap by reviewing known explicit and implicit set-theoretic formulaes on the real line for the deterministic cases, classifying them by structural features, and highlighting avenues for further investigation.
1.4 Outline of the Paper
This survey focuses exclusively on deterministic Cantor sets on the real line, examined through a set-theoretic lens, and highlights several descriptive perspectives that remain undiscovered. Section 3 reviews the necessary background in descriptive set theory: the Borel hierarchy is recalled, and four nested levels of set-theoretic representation—ranging from the broadest class of Cantor sets to the narrowest class of construction—are delineated. Section 4 then supplies descriptive formulas for three families of Cantor sets; the middle-third Cantor set lies in the intersection of all three, thereby serving as a unifying example. Section 5 enumerates current challenges, states open problems, and sketches directions for future work. An appendix collects intermediate background, and interim results, and provides annotated computer code that automates some of the constructions, offering a platform for subsequent mathematical investigations.
2 Set‑Theoretic Foundations
This section deals with the set theoretic foundations for description of Cantor sets. We start with fundamental definitions of $F_{\sigma},$ $G_{\delta}$ and the Borel Hierarchy [53, 54]. Then, we present the proposed set theoretic representations of Cantor sets in hierarchy from macro-level to micro-level in four levels: (i) The general representation, (ii) the nested representation, (iii) the Iterated Function System representation, and, (iv) the $q$- ary expansion.
2.1 The $F_{\sigma}$ and $G_{\delta}$ Properties and the Borel Hierarchy
Let $\mathcal{B}([0,1])$ denote the Borel $\sigma$-algebra on the closed unit interval. Set
$\Sigma^{0}_{1}:=\{\,U\subseteq[0,1]\mid U\text{ is open}\},\qquad\Pi^{0}_{1}:=\{\,F\subseteq[0,1]\mid F^{c}\in\Sigma^{0}_{1}\}\;(=\text{ all closed sets}),$
and define the Borel hierarchy inductively for $n\geq 1$ by:
$\Sigma^{0}_{n+1}:=\Bigl\{\bigcup_{k=1}^{\infty}A_{k}\,\Bigm|\,A_{k}\in\Pi^{0}_{n}\Bigr\},\qquad\Pi^{0}_{n+1}:=\Bigl\{\bigcap_{k=1}^{\infty}B_{k}\,\Bigm|\,B_{k}\in\Sigma^{0}_{n}\Bigr\}.$
The classes $\Sigma^{0}_{2}$ and $\Pi^{0}_{2}$ are traditionally denoted:
[…]
Thus, every closed subset of $[0,1]$ is $F_{\sigma}$, and every open subset of $[0,1]$ is $G_{\delta}$. Also, the middle-third Cantor set is simultaneously $F_{\sigma}$ and $G_{\delta}$ [54].
2.2 General Representation
The most general set-theoretic representation of Cantor sets is introduced in 1999 [55]. To begin with, let $I_{0}=[0,1]$ be the closed unit interval and let $\{I_{n}^{*}\}_{n\geq 1}$ be a (finite or countable) family of pairwise–disjoint open sub‑intervals contained in $I_{0}$. The general definition of Cantor set $\Gamma$ is given by:
$\Gamma=I_{0}-\bigcup_{n=1}^{+\infty}I_{n}^{*}.$ (1)
Considering closed sets $I_{n}=I_{0}-I_{n}^{*}(n\in\mathbb{N})$ we have the following equivalent for the most general representation of Cantor set:
$\Gamma=\bigcap_{n=0}^{+\infty}I_{n}.$ (2)
This definition covers many non-fractal subsets of the closed unit interval. Furthermore, by definition $\Gamma$ is closed subset of $[0,1]$ and hence a $F_{\sigma}$ set. Also, given that for the Euclidean distance $d_{E}$ we have $\Gamma=\cap_{n=1}^{+\infty}\Gamma^{(n)}$ where $\Gamma^{(n)}:=\{x\in[0,1]|d_{E}(x,\Gamma)<\frac{1}{n}\}(n\in\mathbb{N})$ are open sets, it follows that $\Gamma$ is a $G_{\delta}$ set.
2.3 Nested Representation
A special case of the general representation of Cantor sets is the nested representation introduced in 2014 [4] where a nested structure is imposed on their construction process as follows. Let $\bigl(s_{n}\bigr)_{n\geq 0}$ be a strictly increasing sequence of natural numbers with $s_{0}=1$ and $s_{n}\to\infty$. Construct closed sets $I_{n}\subset[0,1]$ for given $n\geq 0$ recursively as follows:
  1. 1.
    Base step. Set $I_{0}=[0,1]$ and denote its single component by $I_{0,1}$.
  2. 2.
    Inductive step. Assume $I_{\,n-1}=\bigcup_{k=1}^{s_{\,n-1}}I_{\,n-1,k}$ is the disjoint union of $s_{\,n-1}$ closed intervals. Subdivide each $I_{\,n-1,k}$ into $m_{n,k}\geq 1$ pairwise–disjoint closed sub‑intervals. From the entire collection of these sub‑intervals choose exactly $s_{n}$ of them (at least one descendant from every parent), relabel them $I_{\,n,1},\dots,I_{\,n,s_{n}}$, and set:
    $I_{n}$ $=$ $\bigcup_{k=1}^{s_{n}}I_{\,n,k}\subseteq I_{\,n-1}\ \ (n\in\mathbb{N}).$ (3)
Assume moreover that the intervals shrink, i.e. $\max_{1\leq k\leq s_{n}}\lvert I_{\,n,k}\rvert\;\xrightarrow[]{}\;0.$ Then $I_{0}\supset I_{1}\supset I_{2}\supset\cdots,\text{with}\ I_{n}=\bigcup_{k=1}^{s_{n}}I_{\,n,k}\ \text{and}\ \ s_{n}\uparrow+\infty.$ The associated Cantor‑set limit set is:
$\Gamma$ $=$ $\bigcap_{n=0}^{+\infty}I_{n}.$ (4)
This definition encompasses a broad class of real-line fractals, characterized by their symmetry, Lebesgue measure, and key topological attributes—most notably perfectness and nowhere denseness.
2.4 Iterated Function System (IFS) Representation
A special instance of the nested representation of Cantor sets is obtained by viewing them as attractors of an iterated-function system (IFS) of affine contractions [4]. In this case, at each stage of the construction the surviving sub-intervals arise as images of the unit interval under a fixed finite family of affine maps. To see construction, we recall the relevant terminology as follows. First, we consider the following special maps:
  1. 1.
    A map $T:[0,1]\to[0,1]$ is affine if $T(tx_{1}+(1-t)x_{2})=tT(x_{1})+(1-t)T(x_{2})$ for all $t,x_{1},x_{2}\in[0,1]$.
  2. 2.
    It is a contraction if $\lvert T(x_{1})-T(x_{2})\rvert\leq r\lvert x_{1}-x_{2}\rvert$ for some $r\in(0,1)$ and all $x_{1},x_{2}\in[0,1]$.
Second, every affine contraction can be written $T(x)=rx+s$ with $r,s\in[0,1]$ and $r<1$ [2]; we call $r=\operatorname{cont.coeff}(T)$ its contraction coefficient. Third, let $\mathcal{T}=\{T_{k}\}_{k=1}^{K}\;(K\geq 2)$ be a finite family of affine contractions with $\max_{1\leq k\leq K}\operatorname{cont.coeff}(T_{k})<1$. Such a family constitutes an IFS. For a non-empty compact set $I\subset[0,1]$ define:
$T(I)$ $=$ $\bigcup_{k=1}^{K}T_{k}(I).$ (5)
The operator $T$ is itself a contraction with ratio $r=\max_{1\leq k\leq K}\operatorname{cont.coeff}(T_{k})$ [4]. Hence the sequence $\{T^{\circ n}([0,1])\}_{n=0}^{+\infty}$ where $T^{\circ n}$ denotes the $n$-fold composition and $T^{\circ 0}=\mathrm{id}$ converges and we may set the Cantor set as:
$\Gamma$ $=$ $\bigcap_{n=0}^{+\infty}T^{\circ n}([0,1]).$ (6)
Here, each $T^{\circ n}([0,1])$ is the disjoint union of $K^{n}$ closed intervals of length at most $r^{\,n}$, so their diameters tend to 0 as $n\longrightarrow+\infty.$ Furthermore, the Cantor set satisfies the invariance relation:
$\Gamma$ $=$ $T(\Gamma)\;=\;\bigcup_{k=1}^{K}T_{k}(\Gamma).$ (7)
Thus, the affine–IFS description furnishes the more specific and informative representation in the nested hierarchy of set-theoretic representations of Cantor sets.
2.5 q-Ary Expansion
The $q$-Ary expansion of Cantor sets is indeed an special case of their IFS representation when the involved affine maps have equal slopes. In details, let $q\in\mathbb{N}+2,$ and $A\subset\{0,1,\cdots,q-1\},|A|=K(2\leq K\leq q).$ For every digit $a\in A$ define the $a^{th}$ strict contraction $T_{a}$ with slope $\frac{1}{q}$ and y-intercept $\frac{a}{q}:$
$T_{a}$ $:$ $[0,1]\longrightarrow[0,1]$
$T_{a}(x)$ $=$ $\frac{1}{q}x+\frac{a}{q}.$ (8)
Now, the IFS representation of the Cantor set $\Gamma$ is given by:
$\Gamma$ $=$ $\bigcap_{n=0}^{+\infty}I_{n}:I_{0}=[0,1],I_{n}=\bigcup_{a\in A}T_{a}(I_{n-1}),\ \ (n\in\mathbb{N}).$ (9)
This special case has the following properties:
  1. 1.
    Each component of $I_{n}$ is a closed interval of length $q^{-n}.$
  2. 2.
    There are $K^{n}$ such components with one for each finite digit block $\vec{a}=(a_{1},\cdots,a_{n}):a_{k}\in A(1\leq k\leq n).$
  3. 3.
    The system is nested: $I_{n}\supset I_{n-1}(n\in\mathbb{N}).$
A straightforward verification by mathematical induction shows that:
$I_{n}$ $=$ $T_{a_{1}}\circ\cdots\circ T_{a_{n}}([0,1])=[\sum_{k=1}^{n}\frac{a_{k}}{q^{k}},\sum_{k=1}^{n}\frac{a_{k}}{q^{k}}+\frac{1}{q^{n}}],\ \ (n\in\mathbb{N}).$ (10)
Also, the map $F_{n}=T_{a_{1}}\circ\cdots\circ T_{a_{n}}$ is $q^{-n}$-Lipschitz continuous $(n\in\mathbb{N})$ [56]. Consequently, $x\in\Gamma$ if and only if there is a unique (up to the standard endpoint ambiguity) infinite digit string $\vec{a}^{*}=(a_{1},a_{2},a_{3},\cdots)\in A^{\mathbb{N}}$ such that $x=\sum_{n=1}^{+\infty}\frac{a_{n}}{q^{n}}.$ Hence, we have the $q$- ary expansion of the Cantor set:
$\Gamma$ $=$ $\{x\in[0,1]|x=\sum_{n=1}^{+\infty}\frac{a_{k}}{q^{n}},a_{n}\in A\}.$ (11)
Notation: Henceforth, we denote $\Gamma(\alpha,K)$ middle $\alpha$- Cantor set defined by $K$ affine maps.
In the upcoming subsections, we study, explicit and recursive representation of three families of Cantor sets given their associated $q$-ary expansions.
3 Descriptive Formulas for Cantor Sets
3.1 Thin Family #1
The earliest set-theoretic explicit description of Cantor sets appears in 2006 [57], where the author treats the symmetric thin family generated by at least two affine maps in an iterated-function system. This construction yields a fully constructive proof of the Hausdorff-Dimension Theorem, thereby establishing the existence of the corresponding fractals for any given Hausdorff dimension. Subsequent work extends the argument to broader deterministic and stochastic settings [47, 48] .
Theorem 3.1
Let $\Gamma_{1}(\frac{1}{q},\frac{q+1}{2}):=\{x\in[0,1]|x=\sum_{n=1}^{+\infty}\frac{a_{n}}{q^{n}}\ ,\ a_{n}=0,2,4,\cdots,q-1\}$ where $q\in 2\mathbb{N}+1.$ Then:
$\Gamma_{1}(\frac{1}{q},\frac{q+1}{2})$ $=$ $[0,1]-\bigcup_{n=1}^{+\infty}\bigcup_{k=0}^{q^{n-1}-1}\bigcup_{r=1}^{\frac{q-1}{2}}(\frac{qk+(2r-1)}{q^{n}},\frac{qk+2r}{q^{n}}):\ \ (q\in 2\mathbb{N}+1).$ (12)
In particular, for the case $q=3$ we obtain the explicit formulae for the middle-third Cantor set $C=\Gamma_{1}(\frac{1}{3},2)$ as in the second formulae in Corollary 3.4.
3.2 Thin Family #2
In the same year, the explicit set-theoretic description of a second family of Cantor sets was presented [58]. As with the earlier construction, the sets obtained are thin and symmetric; however, their underlying iterated-function system involves only two affine contractions, in contrast to generally larger number used for the first family.
Theorem 3.2
Let $\Gamma_{2}(\frac{q-2}{q},2):=\{x\in[0,1]|x=\sum_{n=1}^{+\infty}\frac{a_{n}}{q^{n}}\ ,\ a_{n}=0,q-1\}$ where $q\in\mathbb{N}+2.$ Then:
$\Gamma_{2}(\frac{q-2}{q},2)$ $=$ $[0,1]-\bigcup_{n=1}^{+\infty}\bigcup_{k=0}^{q^{n-1}-1}(\frac{qk+1}{q^{n}},\frac{qk+2}{q^{n}}):\ \ (q\in\mathbb{N}+2).$ (13)
Similar to the Thin Family $\Gamma_{1}$, for the case $q=3$ we obtain the explicit formulae for the middle-third Cantor set $C=\Gamma_{2}(\frac{1}{3},2)$ as in the second formulae in Corollary 3.4.
3.3 Augmented Thick Family #1
In both preceding cases we provided closed-form descriptions for two families of measure-zero Cantor sets. This naturally raises the question: Does an analogous explicit representation exist for Cantor sets of positive Lebesgue measure? Current evidence suggests a negative answer; no such closed formula is known at the time of writing. Nevertheless, a constant-recursive representation remains feasible [59], and the requisite background together with interim results is presented in Appendix A for interested readers.
Theorem 3.3
Let $\Gamma_{3}(\frac{p}{q},2):=\{x\in[0,1]|x=\sum_{n=1}^{+\infty}\frac{a_{n}}{(2q)^{n}}\ ,\ a_{n}=0,1,\cdots,q-p-1,q+p,\cdots,2q-1,\}$ where $p,q\in\mathbb{N},(p,q)=1,0<\frac{p}{q}\leq\frac{1}{3}.$ Then:
$\Gamma_{3}(\frac{p}{q},2)$ $=$ $\bigcap_{n=0}^{+\infty}\bigcup_{k=1}^{2^{n}}[a_{n,k}(\frac{p}{q}),b_{n,k}(\frac{p}{q})]:$
$a_{0,1}(\frac{p}{q})=0;$
$a_{n,k}(\frac{p}{q})=1_{odd}(k)\Big(a_{n-1,\frac{k+1}{2}}(\frac{p}{q})\Big)+1_{even}(k)\Big(a_{n-1,\frac{k}{2}}(\frac{p}{q})+\frac{(1-3\frac{p}{q})+(1-\frac{p}{q})(2\frac{p}{q})^{n}}{(1-2\frac{p}{q})2^{n}}\Big);$
$b_{0,1}(\frac{p}{q})=1;$
$b_{n,k}(\frac{p}{q})=1_{odd}(k)\Big(b_{n-1,\frac{k+1}{2}}(\frac{p}{q})-\frac{(1-3\frac{p}{q})+(1-\frac{p}{q})(2\frac{p}{q})^{n}}{(1-2\frac{p}{q})2^{n}}\Big)+1_{even}(k)\Big(b_{n-1,\frac{k}{2}}(\frac{p}{q})\Big):$
$1\leq k\leq 2^{n},n\in\mathbb{N}.$
The Computer software code to calculate the end points $a_{n,k}(.),b_{n,k}(.)$ are given in Appendix B with associated examples for $\alpha_{1}=\frac{1}{3},$ and $\alpha_{2}=\frac{1}{4}$ [60]. It is noteworthy to mention that the middle-third Cantor set belongs to all three discussed families. Figure 1 summarizes our above discussions in terms of hierarchy of descriptive representations:
Figure 1: Diagram of descriptive formulas for the Cantor sets from set-theoretic perspective: C refers to the middle-third Cantor set.
Finally, for the case $p=1,q=3$ we obtain the first order recursive formulae for the middle-third Cantor set $C=\Gamma_{3}(\frac{1}{3},2)$ as in the third formulae in the following Corollary 3.4:
Corollary 3.4
The middle-third Cantor set has three equivalent set theoretic descriptive representations:
$C$ $=$ $\{x\in[0,1]|x=\sum_{n=1}^{+\infty}\frac{a_{n}}{3^{n}}\ ,\ a_{n}=0,2\}$ (15)
$=$ $[0,1]-\bigcup_{n=1}^{+\infty}\bigcup_{k=0}^{3^{n-1}-1}(\frac{3k+1}{3^{n}},\frac{3k+2}{3^{n}})$ (16)
$=$ $\bigcap_{n=0}^{+\infty}\bigcup_{k=1}^{2^{n}}[a_{n,k},b_{n,k}]:$
$a_{0,1}=0;a_{n,k}=1_{odd}(k)\Big(a_{n-1,\frac{k+1}{2}}\Big)+1_{even}(k)\Big(a_{n-1,\frac{k}{2}}+\frac{2}{3^{n}}\Big);$
$b_{0,1}=1;b_{n,k}=1_{odd}(k)\Big(b_{n-1,\frac{k+1}{2}}-\frac{2}{3^{n}}\Big)+1_{even}(k)\Big(b_{n-1,\frac{k}{2}}\Big);$
$1\leq k\leq 2^{n},n\in\mathbb{N}.$
4 Discussion
4.1 Summary and Contributions
This survey systematizes four complementary descriptions of Cantor sets—general, nested, IFS, and $q$-ary—and develops explicit formulas for the thin families $\Gamma_{1},\Gamma_{2}$ and a recursive, algorithmic description of the augmented thick family $\Gamma_{3}$ via the endpoint maps $a_{n,k},b_{n,k}$ (for rational $\alpha=p/q\leq 1/3$). The middle–third Cantor set appears as a transparent special case within this unified framework.
Main contributions. This survey has five main contributions to the current literature as follows:
  1. (MC1)
    Unification. We give a single language that connects set-theoretic, nested-interval, IFS, and $q$-ary viewpoints, making it straightforward to translate statements and proofs across representations.
  2. (MC2)
    Explicitness beyond the classical case. For the thin families $\Gamma_{1},\Gamma_{2}$ we provide closed-form descriptions; for the augmented thick family $\Gamma_{3}$ we provide a fully explicit recursion for all retained-interval endpoints $(a_{n,k},b_{n,k})$, enabling exact enumeration of gaps and components at any stage.
  3. (MC3)
    Middle–third as a corollary. The classical middle–third set is obtained immediately as a special instance of our formulas, with equivalence across all four descriptions spelled out.
  4. (MC4)
    Algorithmic reproducibility. The accompanying code (Appendix B) implements the endpoint recursion and derived computations, permitting exact generation, visualization, and interrogation (e.g., gap statistics, finite-stage approximants) of the families.
  5. (MC5)
    Bridging to applications. The explicit gap/endpoint control supports routine computation of quantities used in the analysis of intersections and sumsets (e.g., numerical thickness proxies), and simplifies the construction of self-similar probability measures and distribution functions on these sets.
Continuum-theoretic perspective.
Although Cantor sets are totally disconnected, they occur naturally inside many continua as closed, perfect, nowhere-dense subsets. The explicit endpoint recursion for $\Gamma_{3}$ supplies fine control of gap geometry at every stage. This facilitates: (i) explicit embeddings of $\Gamma_{3}(\alpha)$ into standard continua (e.g., as closed subsets of arcs, dendrites, or Peano continua) with prescribed gap patterns; (ii) constructive “fill-in” procedures that yield continua whose complements are modeled by a chosen $\Gamma_{3}(\alpha)$; and (iii) numerical investigation of intersection phenomena (useful in studying when two embedded Cantor-type sets intersect), where computable gap data and thickness proxies are central. In this sense the paper’s formulas act as a bridge, turning qualitative continuum-theory questions into explicit, checkable computations.
Measure-theoretic perspective (a canonical $\Gamma_{3}$ measure).
For each $\Gamma_{3}(\alpha)$ we can canonically equip the set with a Borel probability measure that generalizes the Cantor–Lebesgue measure [61]. At stage $n$ there are $q^{n}$ retained intervals; assign either uniform weights $q^{-n}$ or, more generally, a fixed weight vector $w=(w_{1},\dots,w_{q})$ with $\sum_{i=1}^{n}w_{i}=1$, propagated multiplicatively along the recursion. This produces a self-similar probability measure $\mu_{\alpha,w}$ supported on $\Gamma_{3}(\alpha)$ whose cumulative distribution function is a Devil’s–staircase generalization of the Cantor function: it is continuous, nondecreasing, constant on deleted gaps, and increases exactly on $\Gamma_{3}(\alpha)$. The explicit endpoints $(a_{n,k},b_{n,k})$ allow fast, exact evaluation of cylinder-set masses and numerical plots of the associated distribution function. This construction is immediate from the framework developed here and is implemented by minor extensions of Appendix B.
Outlook.
Because the geometry of $\Gamma_{3}$ is explicit at every finite stage, the framework supports systematic computation of gap-length statistics, empirical thickness proxies, and intersection behavior under translation—tools that are relevant both to continuum-theoretic embeddings and to questions in additive/fractal geometry. We conclude with open problems on asymmetry, irrational endpoints, variable-gap deletions, the role of symmetry, and common intersections; these problems are now posed in a setting where algorithmic experiments and exact finite-stage calculations are straightforward.
4.2 Future Work and Open Problems
The preceding survey has shown that three natural families of deterministic Cantor sets—two of measure zero with explicit descriptions and one of non–negative measure defined recursively—share a common symmetric architecture and possess rational construction end-points; indeed, the classical middle-third Cantor set lies in the intersection of all three. These parallel features highlight several directions in which the present set-theoretic framework remains incomplete requiring more research by interested mathematicians:
  1. (OP1)
    Asymmetric constructions. Devise explicit or recursive set-theoretic formulas for Cantor sets generated by iterated-function systems (IFS) with $K\geq 2$ maps that do not preserve bilateral symmetry.
  2. (OP2)
    Irrational end-points. Find descriptive formulas when the retained intervals have irrational end-points, again allowing $K\geq 2$ IFS components.
  3. (OP3)
    Variable-gap deletions. Given a sequence $\bigl(\alpha_{n}\bigr)_{n\geq 1}$ with $0<\alpha_{n}<1$ satisfying $1-\sum_{n=1}^{\infty}2^{\,n-1}\alpha_{n}\;\geq 0,$ provide a recursive descriptive formulae for the Cantor set obtained by removing at stage $n$ the open middle interval of length $\alpha_{n}$.
  4. (OP4)
    Role of symmetry. Determine how the solution of problem (iii) varies when symmetry is either enforced or relaxed.
  5. (OP5)
    Common ground. Establish conditions under which the families arising in problems (i)–(iv) admit a non-empty mutual intersection.
Resolving these questions would extend the taxonomy of Cantor sets beyond the symmetric, rational cases treated here and clarify how set-theoretic representations interact with geometric asymmetry, irrational numbers, and variable deletion schemes.
Conclusion
This survey brings four set-theoretic perspectives—general, nested, IFS, and $q$-ary—under one framework and demonstrates how they meet in three representative Cantor-set families containing the middle-third classic. The synthesis clarifies which descriptive formulas are now complete and which remain missing, especially for asymmetric, irrational-endpoint, and variable-gap constructions. By distilling these gaps into a concise list of open problems, the paper sets a clear agenda for future work in descriptive set theory and fractal analysis on the real line.
Funding
This research received no external funding.
Institutional Review Board Statement
Not Applicable.
Informed Consent Statement
Not Applicable.
Data Availability Statement
Not Applicable.
Acknowledgments
Not Applicable.
Conflicts of Interest
The author declares no conflicts of interest.
Abbreviations
The following abbreviations are used in this manuscript:
IFS: Iteration Function System; MC: Main Contribution; OP: Open Problem; SVC: Smith Volterra Cantor
Appendix
Appendix.A. Interim Results
Let $\Gamma_{3}(\alpha,2)$ where $0<\alpha\leq\frac{1}{3}$ be the middle $\alpha$ Cantor set defined by two affine maps. Here, its nested representation is given by $\Gamma_{3}(\alpha,2)=\cap_{n=0}^{+\infty}I_{n}(\alpha)$ where $I_{0}(\alpha)=[0,1]$ and $I_{n}(\alpha)$ is defined recursively by removing $2^{n-1}$ middle open intervals each of the length $\alpha^{n}$ from $I_{n-1}(\alpha)$ $(n\in\mathbb{N}).$ We have the following interim results:
  1. 1.
    Family Name By construction, the Lebesgue measure is given by $\lambda(\Gamma_{3}(\alpha,2))=1-\sum_{n=1}^{+\infty}2^{n-1}\alpha^{n}=\frac{1-3\alpha}{1-2\alpha}(0<\alpha\leq\frac{1}{3}).$ Hence, we augment the "Thick Family" with $0<\alpha<\frac{1}{3}$ with the "Thin case" $\alpha=\frac{1}{3};$ and, call it the "Augmented Thick Family" with $0<\alpha\leq\frac{1}{3}.$
  2. 2.
    2q-Ary Expansion When $0<\alpha=\frac{p}{q}\leq\frac{1}{3},p,q\in\mathbb{N},(p,q)=1$, by above construction, the equality $\frac{p}{q}=\frac{2p}{2q}$, and mathematical induction, a straightforward verification shows the following 2q-Ary expansion:
    $\Gamma_{3}(\alpha,2)=\{x\in[0,1]|x=\sum_{n=1}^{+\infty}\frac{a_{n}}{(2q)^{n}}\ ,\ a_{n}=0,1,\cdots,q-p-1,q+p,\cdots,2q-1\}.$
    This representation facilitates the n-Aray expansion for cases with even $q.$ Furthermore, the middle-third Cantor set has the secondary 6-Ary expansion as $C=\{x\in[0,1]|x=\sum_{n=1}^{+\infty}\frac{a_{n}}{6^{n}}\ ,\ a_{n}=0,1,4,5\}.$
  3. 3.
    Proof of Theorem 3.3 We prove the Theorem in steps (a)-(e).
    1. (a)
      Notation Given above recursive construction it follows that, $\Gamma_{3}(\alpha,2)=\cap_{n=0}^{+\infty}I_{n}(\alpha)$ where $I_{n}(\alpha)=\cup_{k=1}^{2^{n}}I_{n,k}(\alpha)$ and $I_{n,k}(\alpha)=[a_{n,k}(\alpha),b_{n,k}(\alpha)],(1\leq k\leq 2^{n},n\in\mathbb{N}_{0}).$
    2. (b)
      Lemma 4.1
      $\lambda(I_{n,k}(\alpha))=\frac{(2\alpha)^{n+1}+2(1-3\alpha)}{2^{n+1}(1-2\alpha)}:\ \ (1\leq k\leq 2^{n},n\in\mathbb{N}_{0}).$
      Proof. The case for $n=0$ is trivial. Let $n\in\mathbb{N}$ and $J_{n,k}(\alpha)$ be the middle open interval of length $\alpha^{n}$ removed from $I_{n-1,k}(\alpha)$ in the nested construction process i.e. $I_{n-1,k}(\alpha)=I_{n,2k-1}(\alpha)\cup J_{n,k}(\alpha)\cup I_{n,2k}(\alpha).$ Then, given, $\lambda(I_{n,2k-1}(\alpha))=\lambda(I_{n,2k}(\alpha))$ we have: $\lambda(I_{n-1,k}(\alpha))=2\lambda(I_{n,k}(\alpha))+\alpha^{n}.$ Consequently, we have a first order recursive relation:
      $\lambda(I_{n,k}(\alpha))=\frac{1}{2}\lambda(I_{n-1,k}(\alpha))-\frac{\alpha^{n}}{2},\lambda(I_{0,1}(\alpha))=1.$
      Finally, the assertion follows by mathematical induction on $n\in\mathbb{N}$. $\Box$
    3. (c)
      Lemma 4.2
      Let $\Delta_{n,k}(\alpha)=\lambda(I_{n,k}(\alpha))+\alpha^{n}\ \ (1\leq k\leq 2^{n},n\in\mathbb{N}_{0}).$ Then,
      $\Delta_{n,k}(\alpha)=\frac{(1-3\alpha)+(1-\alpha)(2\alpha)^{n}}{(1-2\alpha)2^{n}}\ \ (1\leq k\leq 2^{n},n\in\mathbb{N}_{0}).$
      Proof. This is straightforward result from Lemma 4.1. $\Box$
    4. (d)
      Lemma 4.3
      Let $a_{n,k}(\alpha),b_{n,k}(\alpha),(1\leq k\leq 2^{n},n\in\mathbb{N}_{0})$ be defined as in above recursive construction. Then:
      $a_{0,1}(\alpha)$ $=$ $0,$
      $a_{n,k}(\alpha)$ $=$ $1_{odd}(k)\Big(a_{n-1,\frac{k+1}{2}}(\alpha)\Big)+1_{even}(k)\Big(a_{n-1,\frac{k}{2}}(\alpha)+\Delta_{n,k}(\alpha)\Big),$
      $b_{0,1}(\alpha)$ $=$ $1,$
      $b_{n,k}(\alpha)$ $=$ $1_{odd}(k)\Big(b_{n-1,\frac{k+1}{2}}(\alpha)-\Delta_{n,k}(\alpha)\Big)+1_{even}(k)\Big(b_{n-1,\frac{k}{2}}(\alpha)\Big),$
      $(1\leq k\leq 2^{n},n\in\mathbb{N}_{0}).$
      Proof. This is straightforward result from definition of $\Delta_{n,k}(\alpha)$ in Lemma 4.2, and comparing the end points of closed intervals in the following equations:
      $[a_{n-1,k}(\alpha),b_{n-1,k}(\alpha)]$ $=$ $I_{n-1,k}(\alpha)=I_{n,2k-1}(\alpha)\cup J_{n,k}(\alpha)\cup I_{n,2k}(\alpha)$
      $=$ $[a_{n,2k-1}(\alpha),b_{n,2k-1}(\alpha)]\cup(b_{n,2k-1}(\alpha),a_{n,2k}(\alpha))\cup[a_{n,2k}(\alpha),b_{n,2k}(\alpha)],$
      $(1\leq k\leq 2^{n},n\in\mathbb{N}_{0}).$
      $\Box$
    5. (e)
      Completion Finally, by considering $0<\alpha=\frac{p}{q}\leq\frac{1}{3}$ in Lemma 4.2 and substituting corresponding $\Delta_{n,k}$ in the equations presented in Lemma 4.3, the proof for the Theorem 3.3 is completed. $\Box$
Appendix.B. R Software Code
The following function in R software code produces the end points of the intervals of the recursive representation of "Augmented Thick Family" $\Gamma_{3}(\frac{p}{q},2)$ for given pair $(n,k):1\leq k\leq 2^{n},n\in\mathbb{N}.$ It presents two examples of calculations one for the thin Cantor set and another for the thick Cantor set:
[…]
a_nk <- local({
  ## simple memorization (saves repeated work)
  cache <- new.env(parent = emptyenv())
  f <- function(n, k, p, q) {
    if (n == 0L) {                     # base level
      if (k != 1L) stop("For n = 0 you must take k = 1")
      return(0)
    if (k < 1 || k > 2ˆn)
      stop("k must satisfy 1 <= k <= 2ˆn")
    ## build a unique key for memorisation
    key <- paste(n, k, p, q, sep = "-")
    if (exists(key, envir = cache, inherits = FALSE))
      return(cache[[key]])
    r <- p / q                         # the ratio p/q
    ## recursive step
    if (k %% 2L == 1L) {               # k is odd
      val <- f(n - 1L, (k + 1L) / 2L, p, q)
    } else {                           # k is even
      prev <- f(n - 1L, k / 2L, p, q)
      term <- ((1 - 3 * r) + (1 - r) * (2 * r)ˆn) /((1 - 2 * r) * 2ˆn)
      val <- prev + term
    cache[[key]] <- val                # store and return
    val
  f
[…]
b_nk <- local({
  cache <- new.env(parent = emptyenv())  #memorization
  f <- function(n, k, p, q) {
    if (n == 0L) {                  # base case
      if (k != 1L) stop("For n = 0 you must take k = 1")
      return(1)
    if (k < 1L || k > 2Lˆn)
      stop("k must satisfy 1 <= k <= 2ˆn")
    key <- paste(n, k, p, q, sep = "-")
    if (exists(key, envir = cache, inherits = FALSE))
      return(cache[[key]])
    r    <- p / q
    term <- ((1 - 3 * r) + (1 - r) * (2 * r)ˆn) /((1 - 2 * r) * 2ˆn)
    ## recursive step
    if (k %% 2L) {                               # k  is odd
      val <- f(n - 1L, (k + 1L) %/% 2L, p, q) - term
    } else {                                     # k is even
      val <- f(n - 1L, k %/% 2L, p, q)
    cache[[key]] <- val
    val
  f
##  Intervals Endpoints Calculator
EndpointCalculator <- function(n, p, q) {
  ## ---------- base‑level endpoints ----------
  a0 <- a_nk(0L, 1L, p, q)
  b0 <- b_nk(0L, 1L, p, q)
  ## keep the original “too‑large” denominator
  denom <- qˆ(n + 1L)
  ## ---------- level‑n endpoints ----------
  res1 <- sapply(1:2ˆn, function(k) a_nk(n, k, p, q))
  res2 <- sapply(1:2ˆn, function(k) b_nk(n, k, p, q))
  num1 <- c(a0 * denom, round(res1 * denom))
  num2 <- c(b0 * denom, round(res2 * denom))
  ## ---------- LABELS (divide out one factor of q) ----------
  display_denom <- denom / q   #  qˆ(n+1) / q  ->  qˆn
  col_labs       <- paste0("k=", 0:2ˆn)
  fractions1     <- paste0(num1 / q, "/", display_denom)
  fractions2     <- paste0(num2 / q, "/", display_denom)
  ## ---------- NEW ROW ----------
  combo <- paste0("[", fractions1, ",", fractions2, "]")
  out1   <- rbind("a_(n,k)(α)"              = fractions1,
                 "b_(n,k)(α)"              = fractions2,
                 "[a_(n,k)(α),b_(n,k)(α)]" = combo)
  colnames(out1) <- col_labs
  out2<-t(out1)
  return(out2)
##  Examples
## Example (1) : Thin Cantor Set
EndpointCalculator(n=2, p=1, q=3)
#      a_(n,k)(α) b_(n,k)(α) [a_(n,k)(α),b_(n,k)(α)]
# k=0 "0/9"      "9/9"      "[0/9,9/9]"
# k=1 "0/9"      "1/9"      "[0/9,1/9]"
# k=2 "2/9"      "3/9"      "[2/9,3/9]"
# k=3 "6/9"      "7/9"      "[6/9,7/9]"
# k=4 "8/9"      "9/9"      "[8/9,9/9]"
## Example (2) : Thick Cantor Set
EndpointCalculator(n=2, p=1, q=4)
#      a_(n,k)(α) b_(n,k)(α) [a_(n,k)(α),b_(n,k)(α)]
# k=0 "0/16"     "16/16"    "[0/16,16/16]"
# k=1 "0/16"     "2.5/16"   "[0/16,2.5/16]"
# k=2 "3.5/16"   "6/16"     "[3.5/16,6/16]"
# k=3 "10/16"    "12.5/16"  "[10/16,12.5/16]"
# k=4 "13.5/16"  "16/16"    "[13.5/16,16/16]"